OpenStudy (anonymous):

Can somebody please check if my work is correct for the following question: integral (2(xcosx-sinx))/x^2 =2 integral (xcosx-sinx)/x^2 =2 integral (xcosx)/x^2 - (sinx)/x^2 =2 integral cosx/x - integral sinx/x^2 =2 integral 1/x * cosx - x^-2 * sinx =2log(x)sinx-1/x * cosx +c

OpenStudy (solomonzelman):

This is your answer, the last line ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i believe it still can be simplified

OpenStudy (anonymous):

but im not sure

OpenStudy (solomonzelman):

I am not so sure, what you are doing after 2 integral cosx/x - integral sinx/x^2

OpenStudy (anonymous):

2 integral 1/x * cosx - integral x^-2 * sinx

OpenStudy (sidsiddhartha):

2 integral[ 1/x * cosx - x^-2 * sinx] upto this step its alright now from here u should use product rule of integration on integral[[ 1/x * cosx] \[\int\limits_{}^{}uvdx=u \int\limits_{}^{}vdx-\int\limits_{}^{}[\frac{ d }{ dx }(u)*\int\limits_{}^{}vdx]dx\]

OpenStudy (solomonzelman):

it is supposed to be 2 integral 1/x * cosx - 2 integral x^-2 * sinx not just 2 integral 1/x * cosx - integral x^-2 * sinx

OpenStudy (solomonzelman):

2(a-b) is not equal to 2a-b

OpenStudy (sidsiddhartha):

yeah

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ 2\int_{ }^{ } \frac{\cos(x)}{x}- 2\int_{ }^{ } \frac{\sin(x)}{x^2} }\)

OpenStudy (solomonzelman):

then you get \(\Large\color{blue}{ 2\int_{ }^{ } \frac{1}{x}\cos(x)- 2\int_{ }^{ } \frac{1}{x^2}\sin(x) }\)

OpenStudy (anonymous):

ok now i see where i made the mistake. so in other words if i have a constant outside the integral that constant also applies to the second integral?

OpenStudy (sidsiddhartha):

yeah now the first intrgral will be \[(1/x)\int\limits_{}^{}cosxdx-\int\limits_{}^{}[(-1/x^2)(-sinxdx)\] and look they are getting cancelled out

OpenStudy (sidsiddhartha):

*it will be only (sinxdx) not (-sinxdx)

OpenStudy (solomonzelman):

@sidsiddhartha you can not take 1/x out, it is not a constant

OpenStudy (solomonzelman):

you did that in your last black latex integral.

OpenStudy (anonymous):

i agree solomon x's are supposed to be in not out

OpenStudy (solomonzelman):

I would just go with \(\Large\color{blue}{ 2\int_{ }^{ } \frac{\cos(x)}{x}- 2\int_{ }^{ } \frac{\sin(x)}{x^2} }\)

OpenStudy (anonymous):

so it should be 2logx sinx - 2 (x^-1/-1) * -cosx

OpenStudy (sidsiddhartha):

@SolomonZelman i'm not taking i/x out i'm using uv substitution in which 1/x is =u and cosx=v

OpenStudy (sidsiddhartha):

just see the formula i wrote above for uv substitution @SolomonZelman

OpenStudy (anonymous):

2logxsinx-2/xcosx

OpenStudy (anonymous):

would that be the answer?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ 2\int_{ }^{ } \frac{\cos(x)}{x}- 2\int_{ }^{ } \frac{\sin(x)}{x^2} }\) \(\Large\color{blue}{ [2Ci(x)+C]~~- 2(~~Ci(x) - \frac{\sin(x)}{x}~~)~~+C}\) \(\Large\color{blue}{ 2Ci(x)~~- 2(~~Ci(x) - \frac{\sin(x)}{x}~~)~~+C}\) \(\Large\color{blue}{ 2~(~~~Ci(x)-~~Ci(x) - \frac{\sin(x)}{x}~~)~~+C}\) \(\Large\color{blue}{ 2~(- \frac{\sin(x)}{x}~~)~~+C}\)

OpenStudy (solomonzelman):

it should be a positive sin(X) .

OpenStudy (solomonzelman):

2sin(X) ------ + C x

OpenStudy (solomonzelman):

I am not very good at integration... I could have been better if I took calc1, but the highest math course I took is trig -:(

OpenStudy (anonymous):

@SolomonZelman you are pretty good at integration by the looks of it :)

OpenStudy (anonymous):

could you please teach me the steps on how you did it

OpenStudy (solomonzelman):

yes, myininaya told me I was good at it too, when solved an integral in a couple of steps. http://openstudy.com/users/solomonzelman#/updates/53dfebc5e4b0ca0415954e5f

OpenStudy (solomonzelman):

Well... I read some staff in a calc 1 book. About it and practiced some probs.

OpenStudy (anonymous):

so is this step correct? 2 integral (cosx/x - sinx/x^2) dx

OpenStudy (anonymous):

if its correct what is the next step?

OpenStudy (anonymous):

would it be followed by: 2 integral cosx/x - 2 integral sinx/x^2

OpenStudy (sidsiddhartha):

\[2\int\limits_{}^{}cosx/x-2\int\limits_{}^{}sinx/x^2\] \[2[\frac{ 1 }{ x }\int\limits_{}^{}cosxdx-\int\limits_{}^{}(d/dx(1/x)*\int\limits_{}^{}cosxdx)-2\int\limits_{}^{}sinx/x^2\] \[2(1/x)\int\limits_{}^{}cosxdx+2\int\limits_{}^{}(1/x^2)sinxdx-2\int\limits_{}^{}(1/x^2)sinxdx\] so they are getting calcelled out \[=\frac{ 2 }{ x }\int\limits_{}^{}cosxdx=(2/x)sinx=\frac{ 2*sinx }{ x }\] same thing

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (sidsiddhartha):

if u havent learned uv subs then u'll not understand it @Chad123

OpenStudy (anonymous):

i have learnt it but in my class the teacher explained very well that we avoid putting x's outside the integral only constants

OpenStudy (sidsiddhartha):

are u able to grasp it then?

OpenStudy (anonymous):

but you havent used any u's in your explanation thats why it is not clear

OpenStudy (anonymous):

having things on the screen is different when written on paper

OpenStudy (anonymous):

i would like someone to take me through it step by step not write the whole solution

OpenStudy (sidsiddhartha):

its not just UV its LIATE method L=log I=inverse A=algebric T=trigonometric E=exponential i can explain it to u later but i gotta go now sorry

OpenStudy (anonymous):

@SolomonZelman are you there?

OpenStudy (solomonzelman):

Yes, I am now here... my friend is moving, so I had to help out.

OpenStudy (anonymous):

great im stuck :)

OpenStudy (anonymous):

i reposted the question